OpenStudy (anonymous):

Solve 25x^2 = 12(5x-3) thanks :)

OpenStudy (anonymous):

still need this? \[25x^2=60x-36\] \[25x^2-60x+36=0\] \[(5x-6)^2=0\] \[5x=6\] \[x=\frac{6}{5}\]